At December 31, 2017, Albrecht Corporation had outstanding 248,000 shares of common stock and 14,000 shares of 9.5%, $100 par value cumulative, nonconvertible preferred stock. On May 31, 2018, Albrecht sold for cash 12,000 shares of its common stock. No cash dividends were declared for 2018. For the year ended December 31, 2018, Albrecht reported a net loss of $326,000.

Required:

Calculate Albrecht's net loss per share for the year ended December 31, 2018. (Enter your answers in thousands. Negative amounts should be indicated by a minus sign.)

Numerator / Denominator = Net Loss per Share
Answer #1

Net Profit for 2018 = -$326000

(Assumed that the net loss is after provision for dividend at 9.5% on preferred stock.)

Net Loss per Share = -326000 / ((248000*5/12)+(260000*7/12))

=-326000 / 255000 = -$1.278 per share